Given the balanced equation: 3 PbCl2 + Al2(SO4)3 → 3 PbSO4 + 2 AlC13.

How many moles of PbSO4 will be formed when 0.4 mole of A12(SO4)3 is consumed?

A) 0.6 mole C) 1.6 moles B) 1.2 moles D) 2.0 moles
Answer #1

Given :- 3 PbCl2 + 1 Al2(SO4)33 PbSO4 + 2 AlC13

Mole Ratio :- 3 1 3 2

Moles (given) :- 0.4 x ( Let Unknown)

1 / 0.4 = 3 / x

1 x = 3 × 0.4

x = 1.2 moles of PbSO4

Option - B (1.2 moles)